ANNOUNCEMENTS ------------- * RHESSI target for today is AR 9934; target for tomorrow is AR 9934. * JOP104 target for today is AR 9919, 9926 (and 9930); target for tomorrow is somewhere on the west limb, exact pointing TBD. * Station 27 identified and fixed a problem with their exciter. This is believed to be the reason for their poor commanding performance over the last few days. * Commanding times for next week: Monday - Command on D27 all day. Tuesday - Command on D66 until 16 UT then D27. Wedesday - Command on D66 until 16 UT then D27. Thursday - Commanding is good all day. Friday - Command on D66 until 16 UT then D/L only until 19 UT then D16. Saturday - Command on D66 until 15 UT then D/L only until 17:30 UT then D16 until 22 UT then D27. Sunday - Command on D66 until 16 UT then D27. FOT REPORT ---------- Spacecraft Status: Nominal Spacecraft Anomalies: None Accomplished Activities: VIRGO, CELIAS Planned Activities: VIRGO, CELIAS, SWAN NOM HGA Table New Tracking Star Upcoming Operations: None. Ground Anomalies: 128/2310 - Late AOS due to 3kW transmitter problem at D46. (2 min, covered by D16) 129/0504 - TM dropout due to SCP interface down at D46. (23 min, recoverable.) SOLAR STATUS ------------ 05/08 DOY 128 - 05/09 DOY 129 EIT daily On-Line Solar Status and observations: http://umbra.nascom.nasa.gov/eit/plan/log_form20020509.html LASCO daily On-Line Solar Status and observations: http://lasco-www.nrl.navy.mil/observations/2001/Todays_obs.html Web Page for Planning: http://www.bbso.njit.edu/arm INSTRUMENT STATUS ----------------- CDS: Nominal. Engineering, ICAL01 with EIT, JOP136. Tomorrow's science plan is TBD, but a lubrication study is planned for 20-22 UT SUMER: Nominal. JOP104 continues. UVCS: Nominal. Flat field study, 15-hour coronal hole study starting 20 UT at 231 degrees, then mini-synoptic. Tomorrow: Coronal hole study and synoptic. LASCO: Nominal. C2/C3 synoptics. EIT: Nominal. Half-resolution 195 CME watch and synoptic, ICAL today. MDI: Nominal. Full disk magnetograms and dopplergrams for continuous contact. TRACE: Nominal. Observing RHESSI region. Switched to UV this morning.